OpenStudy (anonymous):

I need help completing a problem please. Write an equation in slope-intercept form fro the line (-7,0), (0,-7) So far, I have completed............... y=mx+b m= -7-0/0-7 = -7/-7 = 1 Therefore, now in order to find the equation, I think i should be using y-y1=m(x-x1)....is that correct? If so, would it be like y-0=m(x-(-7) Which would be y=x-7 I do not think i'm doing it completely correct..

OpenStudy (anonymous):

that's not correct your problem come from the slope m= -7-0/0-7 = -7/-7 = 1 <= it's wrong m=(ya-yb)/(xa-xb)=(0-(-7))/(-7-0)=-1 !and not 1

OpenStudy (anonymous):

Your m value is wrong. \[m= {-7-0 \over 0-(-7) } = {-7 \over 7} = -1\]

OpenStudy (anonymous):

Nice, i see where I made that mistake. When I did my problem, I did 0-7 and not 0-(-7); big difference. Thank you so much for catching that. So, would it be completed by putting y-0=m(x-(-7) and equal the 7

OpenStudy (anonymous):

y-0=m(x-(-7) <= that's not correct Once you get the slope you have to solve this equation using one of the point y=mx+b y=-x+b 0=-(-7)+b b=-7 so y=-x-7 Now check your answer with the other point (-7,0) y=-7 and -x-7=-0-7=-7 => correct!